The sum of two numbers is 7. Their product is −170. Find the numbers.
Answer by Boreal(15235)

x+y=7
xy=-170
x=7-y
(7-y)*y =-170
7y-y^2=-170
y^2-7y-170=0
(y-17)(y+10)=0
y=17, y=-10
Pick one root, and the other is x.
x+y=7
xy= -170